Jacob is a member of WCC (an LLC taxed as a partnership). Jacob was allocated $160,000 of business income from WCC for the year. Jacob’s marginal income tax rate is 37 percent. The business allocation is subject to 2.9 percent of self-employment tax and 0.9 percent additional Medicare tax. (Round your intermediate calculations to the nearest whole dollar amount.)

rev: 11_08_2019_CS-189656

b. What is the amount of tax Jacob will owe on the income allocation if the income is qualified business income (QBI) and Jacob qualifies for the full QBI deduction?
Answer #1

QBI deduction allows deduction upto 20 perent of QBI income for income earned from domestic business operated as a sole proprietorship or through a partnership. The QBI is subject to limitations, depending on the taxpayer’s taxable income, that may include the type of trade or business, the W-2 wages paid and the unadjusted basis immediately after acquisition of qualified property held . QBI is the net amount of qualified items of income, gain, deduction and loss from any qualified trade or business, including income from partnership.

Hence,

Total income- 160,000

Less: 20% QBI deduction- (32,000)

Taxable income- 160,000-32,000= 128,000

Tax rate- 2.9+.9= 3.8 percent

Tax liability= 128,000*3.8%= $4,864